HUBBLE SPACE TELESCOPE DAILY REPORT #2457 PERIOD COVERED: 0000Z (UTC) 09/14/99 - 0000Z (UTC) 09/15/99 Daily Status Report as of 258/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.2 Completed Three Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.4 Completed WF/PC-2 7501 (Expansion Parallax Distances to Planetary Nebulae) The WF/PC-2 was used to make parallax distance measurements of the planetary nebula IC418. The largest impediment to our understanding of planetary nebulae is the lack of reliable distances. Acquired by generalized statistical methods in all but a handful of cases, individual PNe distances are often uncertain by factors of two or more. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/MA1 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.6 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.7 Completed WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 1.8 Completed Three Sets of WF/PC-2 8192 (A Snapshot Survey of Probable Nearby Galaxies) The WF/PC-2 was used to make a snapshot survey images of a probable nearby galaxies KK250, NGC-5237 and DWING-2. The census of galaxies in the Local Volume of space is very incomplete, particularly at the faint end. As described in yesterday’s report and HSTAR 7252, the acquisition for the first iteration of this proposal failed, the take data flag remained down, and two observations were affected. Otherwise, the observations were executed as scheduled, and no further anomalies were noted 1.9 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was to be used to make parallel observations of nearby galaxies. As described in yesterday’s report and HSTAR 7252, the acquisition for this proposal failed, the take data flag remained down, and one observation was affected. 1.10 Completed STIS/CCD 7444 (Global Mapping of Molecular Oxygen on Ganymede) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of Jupiter’s largest moon Ganymede in order to map its molecular oxygen. This is to study the absorption bands in the spectrum of Ganymede's trailing hemisphere that result from the presence of condensed molecular oxygen trapped on the surface. This is probably resulting from the bombardment of water ice by the Jovian magnetospheric ions. Near the end of this two-and-one-half hour proposal, the loss of lock described in 2.1 and HSTAR 7253 occurred, taking down the take data flag approximately five minutes before the term exp. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.11 Completed Four Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed Two Sets of FGS 8309 (A Link Between Massive Binary Stars and Non-thermal Radio Emission) The Fine Guidance Sensors were used to make astrometric observations of the massive Wolf-Rayet binary stars WR134 and WR136 and several nearby reference stars. Non-thermal radio emission in Wolf-Rayet stars is explained in terms of synchrotron emission from shocks in the stellar wind. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2/STIS/CCD 8169 (HST-Galileo Io Campaign) Both the WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to make observations of Jupiter’s moon Io. Io plays a central role in the Jovian system, yet the mechanisms that couple its volcanoes, the tenuous atmosphere, the escaping corona and the plasma torus remain elusive. In 1999 there is the unique opportunity to combine HST observations with the Galileo close flybys of Io. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 10 Successful: 10 Scheduled Re-acquisitions: 3 Successful: 3 Per HSTAR 7253, there was a loss of lock beginning at 257/100505Z, near the end of the proposal described in 1.10. 2.2 FHST Updates: Scheduled: 23 Successful: 23 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. ESTR-2 was commanded manually to record at 257/1715Z per ROP RD-4. ESTR-2 was commanded to playback at 257/1948Z per ROP RD-1. The STSI MCE-2 reset at 258/051047Z while the low voltage was on and while outside any SAA interval. ROP NS-12 was utilized at 258/0516Z to reset the STIS flight software error counter. MAMA-2 will be recovered at 259/0012Z, the next time high voltage is turned on via normal SMS commanding. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH